OpenStudy (anonymous):

A manufacturer uses a mold to make a part in the shape of a triangular prism. The dimensions of this part are shown below. The question is What estimate is closest to the volume in cubic millimeters of the part? A 65 B 125 C 249 D 446

OpenStudy (anonymous):

stop copying questions from that test your taking

OpenStudy (anonymous):

i'm not geez thats not even the question is a practice one it doesn't even count for anything god stop being so rude

OpenStudy (sleepyhead314):

Hello @Ashtonsmith12 I saw a post from Twitter leading me to your post, so here I am to help! To find the Volume of a Prism you would multiply the Area of the Base times the Height in this case The Base would be the Triangle and the Area of the Triangle = (1/2)(base)*(height) = (1/2)*(14.4)*(2.1) = ? (we can assume that 14.4 was the base for this because it is the side that is perpendicular (right angle) to the given height in the triangle) Then the Height would be = 8.25 So the Volume would be = (area of the Triangle) * (Height) = [answer] (round to the nearest whole number) I hope this helped! If you any questions/concerns in regards to this question, feel free to ask and tag me :)